OpenStudy (anonymous):

Factor the difference of two cubes. X^3^n-1

OpenStudy (anonymous):



OpenStudy (anonymous):

(x^n)^3-(1)^3

OpenStudy (anonymous):

Now: HINT a^3-b^3 = (a-b) (a^2 + ab + b^2)

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

Not really im still confused

OpenStudy (anonymous):

Just use the formula

OpenStudy (anonymous):

which formula was the answer?

OpenStudy (anonymous):

(x^n)^3-(1)^3 HERE suppose a=x^n and b=1

OpenStudy (anonymous):

And use a^3-b^3 = (a-b) (a^2 + ab + b^2)

OpenStudy (anonymous):

ok so that is the answr

OpenStudy (anonymous):

?

OpenStudy (anonymous):

First try yourself

OpenStudy (anonymous):

(x^n-1) (x^2n + x^n +1)